EGI-InSPIRE:Sa1 2012-10-24

From EGIWiki
Jump to: navigation, search
EGI Inspire Main page


Inspire reports menu: Home SA1 weekly Reports SA1 Task QR Reports NGI QR Reports NGI QR User support Reports



Progress of SA1 issues

(SA1) integration of Albania: the site administrators participated to TF12 training and promised progress in the set up of a grid site (SA1) PQ8 Grid software maintenance and support. Waiting for TCB discussions

Milestones/Deliverables

  • D4.6 Operations Architecture: external review completed
  • D4.7 Operations sustainability: no progress last week

SA1.1 Activity Management

  • meetings:
    • GDB for discussion of suspension policy related to the deployment of unsupported glite 3.1/3.2 services
    • ISGC PC phone call
    • fedSM weekly call
    • COD meeting
    • PRACE/EUDAT/EGI workshop organization meeting
    • EGI Operations meeting, chaired and agenda preparation
    • Fedclouds task force
    • COD/CSIRT meeting to evaluate the status of unsupported middleware monitoring process
    • COD meeting to evaluate status of current activities, and promote new ones: revision of current and future probes, support to emerging NGIs and new site administrators, technical revision of automated certification procedure (in progress)

Main activity of the last 7 days was the coordination of activities around the decommissioning of unsupported software, followup of missing/incomplete probes, procedures for ticketing of sites and related escalation

  • Follow-up of suspension cases for low availability
  • Revision of operations service portfolio following gSLM guidelines (preparatory work for D4.7 deliverable on service sustainability)
  • ops VO membership management
  • DTEAM VO membership management
  • operations ticket handling, handling of various A/R esclations
  • Produced assessment on the VO usage of deployed gLite services
  • Follow up of sites/NGIs with missing security contacts
  • Preparation of OMB agenda

SA1.2 Security

  • Continued to provide the infrastructure to monitor obsolete gLite middleware and participated in several discussions with operations and COD people
  • Chasing sites/tickets related to obsolete gLite middleware
  • Still chasing few sites that have not yet upgraded WMS to address vulnerability #4073
  • TI accreditation process completed, with draft version of RFC 2350 (Expectations for Computer Security Incident Response) submitted. Now awaiting response.
  • Start planning for a security training workshop at ISGC2013
  • Ongoing discussions about IRTF staffing/funding
  • Presentation of HEP federated identity management pilot project at HEPiX meeting
  • Presentation on Grid Security Policies, Procedures and activities at Shonan meeting

SA1.3 Staged rollout

  • UMD 1.9.0 release freeze, includes: IGE SAGA, IGE Gridway, BDII Core, CREAM, GFAL/lcg_utils, WMS, ARC
  • Continuation of SW provisioning towards UMD 2.3.0
  • Grid Operations meeting: 22 October

SA1.3 Integration

  • collection of feedback from GLOBUS TF about site certification procedure
  • planning of workshop on use cases for data management across peer infrastructures

SA1.4 Central tools

SA1.5 Accounting

Worked with a number of sites to publish UserDNs.

NDGF now publishing from SGAS via SSM in production.

Final testing of DGAS with SSM.

SA1.6 Helpdesk

  • revision of ticket escalation procedures in case of tickets without activity
  • finalize the next GGUS release 2012-10-24
  • preparing the first meeting of the GGUS advisory board

SA1.7 Support

  • published newsletter
  • usual tickethandling and dashboard work
  • cod meeting
  • prepare cod f2f

Software Support

Working smoothly, no specific issues to report. 
DMSU tickets flow Oct 14 -- Oct 20
assigned 16
back to tpm 0
reassigned to 3rd level 12
solved 3
open DMSU tickets status
assigned 0
in progress 1
waiting for reply 4
on hold 2

Network Support

  • Installed new LFC catalogue machine on gridsrv6-net.dir.garr.it for IPv6 testing
  • Resumed CREAM CE testing and reporting

SA1.8 Availability and core services

  • revision of EGI.eu OLA (Malgorzata) according to structure of operations service catalog. Organization of meetings with technical providers to agree/discuss about suggested service targets
  • DTEAM VOMRS migration in progress
  • DTEAM VO management duties
  • installed new VOMS instance (from UMD2 release) to test VOMRS migration procedure
  • published final A/R reports for September
  • follow up on A/R recomputation requests
  • discussion of recomputations of A/R statistics fro September for all sites affected by problem:

https://tomtools.cern.ch/jira/browse/SAM-2999


Documentation

  • ongoing work on the operations wiki space
  • ongoing work on EGI OLA

Meetings

  • PRACE/EGI/community meeting on data management (October/beginning of November)